OpenStudy (anonymous):

3x(squared)-8x-16=0 solve using an appropriate technique???

OpenStudy (anonymous):

3x^2-12x+4x-16=0 3x(x-4)+4(x-40=0 (3x+4)(x-4)=0 x=4 x=-3/4